OpenStudy (anonymous):

The Rational numbers include which of the following? positive integers all square roots negative integers fractions π

OpenStudy (anonymous):

fist one

OpenStudy (anonymous):

all except square roots. \[5=\frac{ 5 }{ 1 }\] \[-5=\frac{ -5 }{ 1 }\] \[\frac{ 2 }{ 3 },\frac{ -5 }{ 6 }\] \[\pi=\frac{ 22 }{ 7 }or \frac{ 355 }{ 113 }\]